Logarithamic tables are used define wider things in the universe.Suppose log(n) base e is there e referring to what and suppose if base is 2 that referring to what? How the logaritham system related to binary system? How it converts the wider things into smaller? Can you explain with small example?

Grade:12th Pass

1 Answers

Aman Bansal
11 years ago

Dear Jhansi,

The logarithm of a number is the exponent by which another fixed value, the base, has to be raised to produce that number. For example, the logarithm of 1000 to base 10 is 3, because 1000 is 10 to the power 3:1000 = 10 × 10 × 10 = 103. More generally, if x = by, then y is the logarithm of x to base b, and is written y = logb(x), solog10(1000) = 3.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert